OpenStudy (anonymous):

What is the volume of a square pyramid with a slant of 15 and a side of 24?

Directrix (directrix):

If you know the formula, would you post it here, please?

OpenStudy (anonymous):

Volume = 1/3Bh

Directrix (directrix):

Good. I'm going to upload a diagram because I need to show you something about height. The height of the pyramid and the slant height of the pyramid are not the same thing.

OpenStudy (anonymous):

yeah i knew that...i just dont know how to find the height

Directrix (directrix):

So, we have to find the height of the pyramid using the slant height. Give me a moment to mark up the diagram.

OpenStudy (anonymous):

allright....than i would plug those dimensions in?

Directrix (directrix):

Yes, do you see how I got the 12? Half of the base edge. Plug in the values and crank out the height. Then, the volume falls right out.

OpenStudy (anonymous):

so 12 is the height?

Directrix (directrix):

Yes. Did you see that it is a Pythagorean triple -- a 3-multiple of the (3,4,5) triple. So, 3*3, 3*4 and 3*5.

Directrix (directrix):

Do you want to crank out the volume and post what you get? I'll check it.

OpenStudy (anonymous):

YES PLEASE...I CANNOT GET THIS FOR THE LIFE OF ME

Directrix (directrix):

V = 1/3 B h where B is the area of the square base. V = 1/3 * (24*24) * 9 V = 1/3 * 9 * (24*24) V = 3 * 576 V = 1728 cubic units Let me know what you think.

OpenStudy (anonymous):

why was the height 9?

OpenStudy (anonymous):

THANK YOU SOOO MUCH DUDE

Directrix (directrix):

I answered your question incorrectly. --? so 12 is the height? I should have said "no." 12 is in the floor of the pyramid.

Directrix (directrix):

So, 15^2 = h^2 + 12^2 and h = 9.

OpenStudy (anonymous):

that 12 height was getting to me haha...i kept trying and i cluld not get the right answer...thank you soo much haha